OpenStudy (anonymous):

60 months with no interest if you make $3,000 down payment

OpenStudy (anonymous):

What is your question?

OpenStudy (ganpat):

3000$ per month or annul ??

OpenStudy (anonymous):

A car dealer offers you two deals on a car that cost $16,000. Please calculate the monthly payment, given these two payment options the car dealer is offering. Payment Option 1: You can finance the car for 60 months with no interest if you make a $3000 down payment. Payment option 2: You can finance the car for 72 months ( years) with 1% simple annual interest and no down payment . Which monthly payment is lower? Please explain how you arrived to your answer and showyour work for each equation.

OpenStudy (anonymous):

It doesn't say anything about monthly or annual for the $3000

OpenStudy (ganpat):

part A : 3000 down payment... i.e. on the spot and installments on 13K (16K - 3K) for 60 months.. So per month = 13000/ 60 = 216.67 $...

OpenStudy (ganpat):

for second one, i dont know, the direct formula.. i can give a try.. 72 months = 6 yeas.. Simple interest of 1% .. SO interest on principle amount of 16000 = 160 So total amount to be paid = 16000+ 160 = 16160 for 72 months.. So per month = 16160/ 72 = 224.44$..

OpenStudy (ganpat):

Which monthly payment is lower ----> Part A Please explain how you arrived to your answer ----> u have that in front of u.. and show your work for each equation ----> show the same in ur language...

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

Yes, that is correct for the simple interest. Compound interest gets a bit more complicated though. Also, for the second part, it is over 7 years, not 6. So the monthly payment is $16160/84 months = 192.38. Part A however, is still the better option as you are paying less in the end. Remember, lower monthly payments are not always better as most of the time you will be paying more in the end.